This is such a long question, and it's worth 10 points. I feel like I'm headed in the wrong direction, please help.
The average cost of a new car has
been changing approximately
linearly with time. In 1990, the
average cost was $16,950, and the
average cost was $27,957 in 2008.
a) Let x = the number of years after
1990 and let y = average cost of
a new car. What ordered pairs
can be written to express this
information?

b) Write the equation of the line
through these two points in
point-slope form.

c) Convert the equation of the line
to slope-intercept form and
explain in your own words what
the slope means in the context
of this problem.

d) Based on this linear model, what
will be the average cost of a new
car in 2018?


Answer by rothauserc(4718)

a) (0, 16950), (8, 27957)
b) slope m is (27957 - 16950) / (8 - 0) = 11007 / 8 = 1375.875
use the point (8, 27957)
(y-27957) = 1375.875(x-8)
c) (y-27957) = 1375.875(x-8)
y-27957 = 1375.875x - 11007
y = 1375.875x + 16950
The slope 1375.875 means that for each year after 1990, the averge cost of a new car increases by 1375.875
d) now 2018 - 1990 = 28
y = 1375.875x + 16950
let x = 28, then
y = 1375.875(28) + 16950
y = 38524.5 + 16950
y = 55474.5
therefore the average cost of a new car in 2018 will be $55,474.50
******************************************
makes me wonder how many unsold new cars will be in 2018